# Scalable and Resource-Efficient Second-Order Federated Learning via Over-the-Air Aggregation

### v1

@@ I Introduction
Second-order algorithms were introduced to leverage the Hessian matrix to provide refined update steps, leading to faster convergence. However, despite reducing the number of communication rounds, these methods incur higher computational and communication costs due to the need to compute and transmit the full Hessian matrix. Privacy concerns also arise, as the Hessian contains sensitive information about the local objective function and data. In [6] and [7], the authors address these issues by sharing the inverse Hessian-gradient product vector. However, these approaches fail to scale up as they require computing and storing the full Hessian, leading to computational complexity and storage requirements of $\mathcal{O}(d^{3})$ and $\mathcal{O}(d^{2})$, respectively, where $d$ is the model size. In [8], the inverse of the Hessian is approximated iteratively and recursively multiplied with the gradient to reduce storage needs. Nevertheless, calculating the Hessian-gradient product still entails a complexity of $\mathcal{O}(d^{3})$, making it inefficient for energy-constrained IoT applications. In [9], the authors estimated the Hessian as a diagonal matrix by leveraging the optimizer introduced in [10]. However, the models were updated by performing multiple local iterations before aggregating them, which not only imposed a significant computational burden but also introduced the risk of model drift [11].
@@ III-A Fed-Sophia
The diagonal elements of $\hat{\bm{h}}_{n}^{k}\triangleq\text{diag}(\bm{H}_{n}^{k})$ can be efficiently estimated using the Gauss-Newton-Bartlett (GNB) estimator [10, Algorithm 2]. Using the diagonal elements of the Hessian is a cost-effective and memory-efficient way to retain curvature information across each dimension while also preserving privacy. Hence, the diagonal elements $\hat{\bm{h}}_{n}^{k}$ can be expressed as
$\hat{\bm{h}}_{n}^{k}=B\cdot\hat{\bm{g}}_{n}^{k}\odot\hat{\bm{g}}_{n}^{k},$ (5)
where $\odot$ denotes the element-wise product, $B$ is the mini-batch size, and $\hat{\bm{g}}_{n}^{k}$ is the gradient of the mini-batch loss with labels $\hat{\bm{y}}_{j}$ sampled from the logits function $\bm{\phi}(\bm{\theta}_{n}^{k},\bm{x}_{j})$ on the inputs $\bm{x}_{j}$.
@@
where $\beta_{1}$ and $\beta_{2}$ are tuning parameters. The PS aggregates the received EMA of the gradients and the Hessians as follows
$\bar{\bm{m}}^{k}=\frac{1}{N}\sum_{n=1}^{N}\bm{m}_{n}^{k},\quad\quad\bar{\bm{h}}^{k}=\frac{1}{N}\sum_{n=1}^{N}\bm{h}_{n}^{k}.$ (8)
Using these EMAs, the PS computes the update direction by performing the element-wise division $\bar{\bm{m}}^{k}/\bar{\bm{h}}^{k}$. Employing the update direction $\bar{\bm{m}}^{k}/\bar{\bm{h}}^{k}$ with non-convex functions can result in convergence to a maximum or saddle point, as the Hessian matrix may exhibit rapid changes and inaccuracies, particularly in the early stages of optimization, rendering it unreliable. To address this, we use element-wise clipping to consider values from $\bar{\bm{h}}^{k}$ only when curvature is positive and limit the step size when curvature is steep. Defining the clipping function for a threshold $\gamma>0$ of vector $\bm{z}\in\mathbb{R}^{d}$ as $\widehat{\bm{z}}=\text{clip}(\bm{z},\gamma)$, where $\widehat{z}_{i}=\max(\min(z_{i},\gamma),-\gamma)$, the update in (3) can be written as
@@ III-B Analog Over-The-Air Direction Aggregation Fed-Sophia (OTA Fed-Sophia)
At each iteration $k$, the $N$ clients send their computed EMAs $\bm{m}_{n}^{k}$ and $\bm{h}_{n}^{k}$ through a wireless fading MAC, through $b$ subcarriers over a total of $\tilde{T}\in\{T,2T\}$ time slots, where $b\leq d$ and $T=\lceil\frac{d}{b}\rceil$. To ensure the signals reach the PS simultaneously and utilize the channel’s superposition property, clients perform slot-level synchronization, by using, e.g., timing advance in LTE, which adjusts transmission timing to mitigate misalignment and synchronize signal arrival at the PS [15]. At each time slot $t$ of iteration $k$, each worker $m$ transmits a $b$-length vector through the wireless channel denoted by $\bm{S}_{n}(t)=[S_{n,1}(t),\dots,S_{n,b}(t)]^{T}\in\mathbb{C}^{b}$. The channel output received by the PS over the $i$th subcarrier is given by,
@@
where $h_{n,i}(t)\in\mathbb{C}$ is the fading channel of the $i$th subcarrier between the PS and client $n$, $z_{i}(t)\in\mathbb{C}$ is the additive white Gaussian noise (AWGN) at the PS at time slot $t$ and follows $\mathcal{C}\mathcal{N}(0,1)$. In order to facilitate the updates in (6) and (7), for every element $\bar{\bm{m}}_{i}^{k}$ element in $\bar{\bm{m}}^{k}$ the PS needs to perform
@@
[…]
$\frac{1}{d}\sum_{i=1}^{d}|s_{n,i}(t)|^{2}\leq P_{n},$ (12)
where $P_{n}$ is the maximum power of client $n$. With the assumption that at each time slot $t$, the channel state information (CSI) is known for the PS and the clients, each client $n$ can utilize the knowledge of $h_{n,i}(t)$ and perform the channel inverse operation. However, to avoid the violation of the power constraint imposed on each client we set the transmitted signal $s_{n,i}(t)$ as
@@
Subsequently, the PS divides the expression $\alpha(t)\sum_{n\in N_{i}(t)}\bm{m}_{n,i}^{k}(t)+z_{i}^{k}(t)$ by $\alpha(t)$ and applies matched filtering to the received signal, and the processed received signal becomes $\bar{\bm{m}}_{i}=\sum_{n\in N_{i}(t)}\bm{m}_{n,i}^{k}+\hat{z}_{i}^{k}$, where $\hat{z}_{i}^{k}$ is the matched filtered AWGN. Similarly, after each $\tau$ communication round, the PS obtains the new approximate Hessian $\bm{h}_{n,i}(t)$ and applies matched filtering to construct $\bar{\bm{h}}_{i}^{k}=\sum_{n\in N_{i}(t)}\bm{h}_{n,i}^{k}+\hat{z}_{i}^{k}$. Finally, the parameter server determines the update direction according to (9) and broadcasts the resulting vector to the edge clients via an error-free communication channel to update the local model of the clients. The entire process is provided in Algorithm 1.
@@ IV-A Training Settings
We consider a multi-class classification problem and conduct several experiments to compare the performance of Fed-Sophia with two baselines: FedAvg and DONE using MNIST and CIFAR-10 datasets, respectively. For the simulation settings, the data is distributed among $N=32$ devices in an independent and identically distributed (IID) manner, with each partition split into $75\%$ for training and $25\%$ for testing. We employ two different classification models: an MLP for the MNIST dataset and a CNN for CIFAR-10. The cross-entropy loss function is used for training both models. In the experiments, each communication round involves 1 local iteration for Fed-Sophia and 10 local iterations for FedAvg. For DONE, additional local iterations are necessary due to its convergence dependency on the number of local iterations. For Fed-Sophia and FedAvg, we utilize a batch size of 64, while DONE requires using the full dataset to compute the gradient.
@@ IV-B Communication Network Settings
For OTA Fed-Sophia, analog communication is utilized in the uplink to alleviate the communication bottleneck, while digital communication is employed in the downlink to broadcast the updated model to all clients. Since each element of the update vector is on a subcarrier through a noisy channel on the same subcarrier, all clients need to transmit the entire vector over $\lceil\frac{d}{b}\rceil=\{67,144,352,9350\}$ time slots for the MNIST dataset with an MLP model, Sent140 with an LSTM model, Cifar-10 with a CNN model, and Cifar-100 with a ResNet model [1], respectively. For the digital Fed-Sophia, DONE, and FedAvg, clients transmit the elements of the update directions using $32$-bit representation, and the required number of transmissions becomes $32d$, and therefore the total number of time slots $\tau_{n}$ needed by the $n$th client to transmit the entire update vector to the PS, and the total number of available subcarriers is divided among all $N$ clients, is calculated as follows
@@
The $n$th device’s total energy expenditure throughout the training is composed of parts: (i) the communication component, $E_{t}$, which represents the energy necessary for transmitting and receiving data between the client and the PS, and (ii) the computation component, $E_{c}$, i.e., the energy required to power the hardware, such as CPUs, GPUs, and memory while performing the computation. After $k$ global iterations, the total energy consumed by device $n$ can be expressed as follows
$E_{n,T}(k)\!=\!E_{n,c}(k)\!+\!E_{n,t}(k)\!=\!\sum_{j=1}^{k}e_{n}^{j}\!+\!32d\sum_{j=1}^{k}e_{n,PS}^{j},$ (17)
where $e_{n}^{j}$ is the energy consumed by $n$th client to train and update the model during $j^{\prime}$th communication round, and $e_{n,PS}^{j}$ is the energy needed to transmit one bit to the PS, assuming a 32-bit representation.
@@ IV-C Performance Comparison
In this subsection, we compare the performance of OTA Fed-Sophia against Fed. Sophia, FedAvg, FedProx[16], and DONE in terms of test accuracy on both the MNIST, Sent140, Cifar-10, and CIFAR-100 datasets w.r.t the number of uploads, as computed in (16).
@@ IV-D Computation Time and Energy Comparison
Analyzing the time complexity reveals that FedAvg, FedProx, and (OTA)Fed-Sophia have a linear complexity of $\mathcal{O}(d)$ per client per iteration, while DONE has a cubic complexity of $\mathcal{O}(d^{3})$. Despite their lower time complexities, FedAvg and FedProx require longer time and more energy to achieve the same target accuracy. In Table I, we present the CPU time and emitted carbon footprint needed to reach an $80\%$ accuracy for different algorithms. Our approach outperforms DONE by reaching the target within $85\%$ less CPU time and FedAvg/FedProx by 3×3$\times$3 × less. FedAvg, with the lowest computational complexity, exhibits notably slow convergence, while DONE, despite its high complexity per communication round, requires fewer rounds than FedAvg. However, transmitting updates digitally introduces significant delays, especially for large models. Leveraging the superposition property of wireless channels, OTA Fed-Sophia drastically reduces transmission delays, achieving the target accuracy in one-tenth of the CPU time required by the digital approach. Additionally, Fed-Sophia and OTA Fed-Sophia boast the lowest energy consumption, with DONE consuming the most due to its complex calculations. Overall, Fed-Sophia and OTA Fed-Sophia are the most energy-efficient, consuming only $15\%$ of FedAvg’s energy, $20\%$ of FedProx’s, and $6.7\%$ of DONE’s, with the substantial communication resources OTA Fed-Sophia saves compared to the digital version.
TABLE I: CPU time and carbon footprint required to reach an accuracy of $80\%$ for the different algorithms for the MNIST dataset using an MLP model.

### v3

@@ Abstract
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/LWC.2024.3521027
This work is funded by the European Union under Grant Agreement 101096379. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission (granting authority). Neither the European Union nor the granting authority can be held responsible for them.
@@ I Introduction
Second-order algorithms were introduced to leverage the Hessian matrix to provide refined update steps, leading to faster convergence. However, despite reducing the number of communication rounds, these methods incur higher computational and communication costs due to the need to compute and transmit the full Hessian matrix. Privacy concerns also arise, as the Hessian contains sensitive information about the local objective function and data. In [6], the authors address these issues by sharing the inverse Hessian-gradient product vector. However, these approaches fail to scale up as they require computing and storing the full Hessian, leading to computational complexity and storage requirements of $\mathcal{O}(d^{3})$ and $\mathcal{O}(d^{2})$, respectively, where $d$ is the model size. In [7], the inverse of the Hessian is approximated iteratively and recursively multiplied with the gradient to reduce storage needs. Nevertheless, calculating the Hessian-gradient product still entails a complexity of $\mathcal{O}(d^{3})$, making it inefficient for energy-constrained IoT applications. In [8], the authors estimated the Hessian as a diagonal matrix by leveraging the optimizer introduced in [9]. However, the models were updated by performing multiple local iterations before aggregating them, which not only imposed a significant computational burden but also introduced the risk of model drift [10].
@@ III-A Fed-Sophia
The diagonal elements of $\hat{\bm{h}}_{n}^{k}\triangleq\text{diag}(\bm{H}_{n}^{k})$ can be efficiently estimated using the Gauss-Newton-Bartlett (GNB) estimator [9, Algorithm 2]. Using the diagonal elements of the Hessian is a cost-effective and memory-efficient way to retain curvature information across each dimension while also preserving privacy. Hence, the diagonal elements $\hat{\bm{h}}_{n}^{k}$ can be expressed as $\hat{\bm{h}}_{n}^{k}=B\cdot\hat{\bm{g}}_{n}^{k}\odot\hat{\bm{g}}_{n}^{k}$, where $\odot$ denotes the element-wise product, $B$ is the mini-batch size, and $\hat{\bm{g}}_{n}^{k}$ is the gradient of the mini-batch loss with labels $\hat{\bm{y}}_{j}$ sampled from the logits function $\bm{\phi}(\bm{\theta}_{n}^{k},\bm{x}_{j})$ on the inputs $\bm{x}_{j}$.
@@
where $\beta_{1}$ and $\beta_{2}$ are tuning parameters. The PS aggregates the received EMA of the gradients and the Hessians as $\bar{\bm{m}}^{k}=\frac{1}{N}\sum_{n=1}^{N}\bm{m}_{n}^{k}$, and $\bar{\bm{h}}^{k}=\frac{1}{N}\sum_{n=1}^{N}\bm{h}_{n}^{k}$, respectively. Using these EMAs, the PS computes the update direction by performing the element-wise division $\bar{\bm{m}}^{k}/\bar{\bm{h}}^{k}$. Employing the update direction $\bar{\bm{m}}^{k}/\bar{\bm{h}}^{k}$ with non-convex functions can result in convergence to a maximum or saddle point, as the Hessian matrix may exhibit rapid changes and inaccuracies, particularly in the early stages of optimization, rendering it unreliable. To address this, we use element-wise clipping to consider values from $\bar{\bm{h}}^{k}$ only when curvature is positive and limit the step size when curvature is steep. Defining the clipping function for a threshold $\gamma>0$ of vector $\bm{z}\in\mathbb{R}^{d}$ as $\widehat{\bm{z}}=\text{clip}(\bm{z},\gamma)$, where $\widehat{z}_{i}=\max(\min(z_{i},\gamma),-\gamma)$, the update in (3) can be written as
@@ III-B Analog Over-The-Air Direction Aggregation Fed-Sophia (OTA Fed-Sophia)
At each iteration $k$, the $N$ clients send their computed EMAs $\bm{m}_{n}^{k}$ and $\bm{h}_{n}^{k}$ through a wireless fading MAC, through $b$ subcarriers over a total of $\tilde{T}\in\{T,2T\}$ time slots, where $b\leq d$ and $T=\lceil\frac{d}{b}\rceil$. To ensure the signals reach the PS simultaneously and utilize the channel’s superposition property, clients perform slot-level synchronization, by using, e.g., timing advance in LTE, which adjusts transmission timing to mitigate misalignment and synchronize signal arrival at the PS [14]. At each time slot $t$ of iteration $k$, each client $n$ transmits a $b$-length vector through the wireless channel denoted by $\bm{S}_{n}(t)=[S_{n,1}(t),\dots,S_{n,b}(t)]^{T}\in\mathbb{C}^{b}$. The channel output received by the PS over the $i$th subcarrier is given by,
@@
Subsequently, the PS divides the expression $\alpha(t)\sum_{n\in N_{i}(t)}\bm{m}_{n,i}^{k}(t)+z_{i}^{k}(t)$ by $\alpha(t)$ and applies matched filtering to the received signal, and the processed received signal becomes $\bar{\bm{m}}_{i}=\sum_{n\in N_{i}(t)}\bm{m}_{n,i}^{k}+\hat{z}_{i}^{k}$, where $\hat{z}_{i}^{k}$ is the matched filtered AWGN. Similarly, after each $\tau$ communication round, the PS obtains the new approximate Hessian $\bar{\bm{h}}_{i}^{k}=\sum_{n\in N_{i}(t)}\bm{h}_{n,i}^{k}+\hat{z}_{i}^{k}$. Finally, the parameter server determines the update direction according to (7) and broadcasts the resulting vector to the clients via an error-free communication channel to update the local models. The entire process is provided in Algorithm 1. Although $\bm{m}_{i}^{k}$ is recovered exactly, AWGN introduces an error when $\alpha(t)$ is small, amplifying the noise. Jointly optimizing the transmission power and scaling factor to minimize $\sum_{n=1}^{N}(\bm{m}_{n}^{k}-\bar{\bm{m}}^{k})$ is the right approach, which is left for future work due to its complexity.
@@ IV-A Training Settings
We consider a classification problem and conduct several experiments to compare the performance of OTA Fed-Sophia with three baselines: FedAvg, FedProx[10], and DONE. To demonstrate the performance, we use MNIST, Sent140, CIFAR-10, and CIFAR-100 datasets, employing a multilayer perceptron (MLP), an LSTM, a convolutional NN (CNN), and a ResNet architecture [1], respectively. The data is distributed among $N=32$ clients in an independent and identically distributed (IID) manner, with $75\%$ for training and $25\%$ for testing. The cross-entropy loss function is used to train all models. Each communication round involves 1 local iteration for Fed-Sophia and 10 for FedAvg and FedProx, while DONE requires additional local iterations due to its reliance on multiple updates for convergence. Fed-Sophia, FedAvg, and FedProx use a batch size of 64, whereas DONE processes the full dataset to compute the gradient. To evaluate robustness under non-IID conditions, we also conduct an experiment where data heterogeneity is introduced by limiting each client to a maximum of three labels.
@@ IV-B Communication Network Settings
For OTA Fed-Sophia, analog communication is utilized in the uplink to alleviate the communication bottleneck, while digital communication is employed in the downlink to broadcast the updated model to all clients. Since each element of the update vector is on a subcarrier through a noisy channel on the same subcarrier, all clients need to transmit the entire vector over $\lceil\frac{d}{b}\rceil=\{67,144,352,9350\}$ time slots for the MNIST dataset with an MLP model, Sent140 with an LSTM model, Cifar-10 with a CNN model, and Cifar-100 with a ResNet model, respectively. For the digital Fed-Sophia, DONE, and FedAvg, clients transmit the elements of the update directions using $32$-bit representation, and the required number of transmissions becomes $32d$, and therefore the total number of time slots $\tau_{n}$ needed by the $n$th client to transmit the entire update vector to the PS, and the total number of available subcarriers is divided among all $N$ clients, is calculated as follows
@@
The $n$th device’s total energy expenditure throughout the training is composed of parts: (i) the communication component, $E_{t}$, which represents the energy necessary for transmitting and receiving data between the client and the PS, and (ii) the computation component, $E_{c}$, i.e., the energy required to power the hardware, such as CPUs, GPUs, and memory while performing the computation. After $k$ global iterations, the total energy consumed by device $n$ can be expressed as $E_{n,T}(k)\!=\!E_{n,c}(k)\!+\!E_{n,t}(k)\!=\!\sum_{j=1}^{k}e_{n}^{j}\!+\!32d\sum_{j=1}^{k}e_{n,PS}^{j}$, where $e_{n}^{j}$ is the energy consumed by $n$th client to train and update the model during $j^{\prime}$th communication round, and $e_{n,PS}^{j}$ is the energy needed to transmit one bit to the PS, assuming a 32-bit representation.
@@ IV-C Performance Comparison
In this subsection, we compare the performance of OTA Fed-Sophia against Fed. Sophia, FedAvg, FedProx, and DONE in terms of test accuracy on both the MNIST, Sent140, Cifar-10, and CIFAR-100 datasets w.r.t the number of uploads, as computed in (12).
@@
TABLE I: Comparison of maximum accuracy for IID and non-IID data distributions after $1.5\times 10^{4}$ communication uploads.
OTA Fed-Sophia Fed-Sophia FedAvg DONE FedProx
IID 0.973 0.958 0.890 0.942 0.936
non-IID 0.870 0.858 0.621 0.835 0.862
In the case of non-IID settings OTA Fed-Sophia demonstrates robust performance by maintaining the highest accuracy level as depicted in Table I. In contrast, FedAvg suffers from significant degradation due to model drift, as client updates diverge during local training to minimize communication costs. FedProx mitigates this issue by penalizing deviations from the global model, stabilizing performance in non-IID conditions. OTA FedSophia addresses client deviation by avoiding multiple local updates and compensates for the extra communication overhead through fast convergence and the efficient use of over-the-air aggregation.
@@ IV-D Computation Time and Energy Comparison
FedAvg, FedProx, and (OTA)Fed-Sophia exhibit a linear time complexity of $\mathcal{O}(d)$ per client per iteration, while DONE has a cubic complexity of $\mathcal{O}(d^{3})$. Despite this, FedAvg and FedProx require more time and energy to achieve the same accuracy. (OTA)Fed-Sophia’s efficiency arises from the GNB estimator, which uses gradients to estimate diagonal elements of the Hessian. In Table II, we present the CPU time and emitted carbon footprint needed to reach an $80\%$ accuracy for different algorithms. Our approach outperforms DONE by reaching the target within $85\%$ less CPU time and FedAvg/FedProx by 3×3$\times$3 × less. FedAvg, with the lowest computational complexity, exhibits notably slow convergence, while DONE, despite its high complexity per communication round, requires fewer rounds than FedAvg. However, transmitting updates digitally introduces significant delays, especially for large models. Leveraging the superposition property of wireless channels, OTA Fed-Sophia drastically reduces transmission delays, achieving the target accuracy in one-tenth of the CPU time required by the digital approach. Additionally, Fed-Sophia and OTA Fed-Sophia boast the lowest energy consumption, with DONE consuming the most due to its complex calculations. Overall, Fed-Sophia and OTA Fed-Sophia are the most energy-efficient, consuming only $15\%$ of FedAvg’s energy, $20\%$ of FedProx’s, and $6.7\%$ of DONE’s, with the substantial communication resources OTA Fed-Sophia saves compared to the digital version.
TABLE II: CPU time and carbon footprint required to reach an accuracy of $80\%$ for the different algorithms for the MNIST dataset using an MLP model.